OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). sin 2x - sin 4x = 0

zepdrix (zepdrix):

Are those exponents on the sine functions, or coefficients on the x's?

zepdrix (zepdrix):

\[\large \sin 2x- \sin 4x=0\]Like that? :D

OpenStudy (anonymous):

Yes!

zepdrix (zepdrix):

First of all, we'll need to apply the Sine Double Angle Formula to the second term. Recall this identity:\[\large \sin 2 \theta=2 \sin \theta \cos \theta\] In our problem, we have \(2\theta=4x\), understand how we're going to apply this rule? :D

zepdrix (zepdrix):

\[\large \sin 4x \quad = \quad \sin 2(2x)\]Applying the rule gives us,\[\large 2 \sin(2x) \cos (2x)\] So our problem now becomes,\[\large \sin(2x)-\sin(4x)=0 \\\large \sin(2x)-2\sin(2x)\cos(2x)=0\]

zepdrix (zepdrix):

Bah, sorry if you were typing it out :D haha

zepdrix (zepdrix):

From here, you might notice that each term has a sin(2x) in it. We can factor that out.

zepdrix (zepdrix):

\[\large \sin(2x)\left[1-2\cos(2x)\right]=0\]

OpenStudy (anonymous):

Im so confused...Ive been having a lot of trouble with this lesson

OpenStudy (mertsj):

\[\sin2x-\sin4x=\sin2x-\sin(2x+2x)=\sin2x-[\sin2xcos2x+\cos2xsin2x]=\] \[\sin2x-2\sin2xcos2x=\sin2x(1-2\cos2x)=0\] \[\sin2x=0 or 1-2\cos2x=0\] If sin2x=0, 2x= 0, 180, 360, 540 so x = 0, 90, 180, 270 1-2cos2x=0 cos 2x=1/2 2x= 60, 300, 420, 660 x=30,150, 210, 330

zepdrix (zepdrix):

Mert applied the Angle Sum Identity for Sine, maybe apply that instead of the Sine Double Angle if it makes more sense to you. :D

OpenStudy (anonymous):

Thank you so much, the Angle Sum Identity makes more sense to me than the Sine Double Angle does